OpenStudy (anonymous):

A production company spent 59.5 days filming its first movie, its second movie was filmed in 9 1/4 weeks. find the ration of the first movies filming time to the second movies time. compare in weeks.

OpenStudy (callisto):

59.5 days = 59.5 /7 weeks = 8.5 weeks 9 1/4 weeks = 9.25 weeks First movie: Second movie = 8.5 : 9.25 = 850 : 925 = can you simplify it?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

:)